OpenStudy (anonymous):

A family refers to a groups of two or more people related by birth, marriage, or adoption who reside together. In 2000, in Country A, the average family net worth was $440,000, and there were about 6.9 x 10^7 families. Calculate the total family net worth in Country A in 2000. (Use scientific notation. Round to the nearest tenth as needed.)

OpenStudy (anonymous):

Can someone please help me solve this?

OpenStudy (unklerhaukus):

\[$440,000\times6.9 \times 10^7\]\[=$4.4\times10^5\times6.9 \times 10^7\]\[=$4.4\times6.9 \times10^{5+7}\] and you'll have to divide/multiply by 10 , because scientific notation has only one numeral before the decimal place

OpenStudy (anonymous):

@UnkleRhaukus Thank you for you help thus far, but what do you mean by divide/multiply? Sorry I'm confused.

OpenStudy (unklerhaukus):

do you have the step before that \[=$4.4\times6.9 \times10^{5+7}\] \[=\]

OpenStudy (anonymous):

\[30.36^{12}\] ? i dont know if i am doing this right.

OpenStudy (unklerhaukus):

\[=$4.4\times6.9 \times10^{5+7}\]\[=$30.36\times10^{(5+7)}\]\[=$3.036\times10^{(5+7+1)}\]

OpenStudy (anonymous):

$3 x 10^13

OpenStudy (unklerhaukus):

yeah that is your final solution, did you understand why we had to do the divide/multiply by 10 part?

OpenStudy (anonymous):

I think so, because the number needed to be rounded to the tenth place.

OpenStudy (unklerhaukus):

yeah

OpenStudy (anonymous):

Thank you for your help. I understand it a lot better now.